OpenStudy (anonymous):

The average height of an apple tree is 4.00 meters. How long would it take an apple falling from that height to reach the ground? Given: g = -9.8 meters/second2

OpenStudy (anonymous):

S for distance a for acceleration (in your case g) v0 for starting sped (in this case = 0) S= v0*t + (a*t^2)/2= (a*t^2)/2 a*t^2=2S t^2=(2*S)/a t= square root ->((2*S)/a)